given the diameter of the zinc atom as 2.66 x 10(-8) cm, find the thickness of the zinc coating in the atom?

How many moles of zinc ae in the coating?
1
Expert's answer
2017-10-16T14:47:07-0400
In the question some letters are omitted. May be question is formulated incorrectly. But if suppose that in the problem it is necessary to cover the atom of zinc with zinc atoms. We obtain the problem of Newton numbers [1] . For a three-dimensional space, around the sphere, you can arrange the most dense 12 spheres i.e. 12 atoms of zinc. The thickness of the coating will be equal to the diameter of one atom (i.e. 2.66 x 10-8 cm). Amount of zinc in moles [2] can be calculated using the Avogadro number [3]. Avogadro number is 6.022×1023 mol−1. There are equation: n (amount of substance) = N (number of atoms) /NA (Avogadro constant) = 12 /6.022∙1023= 1,992∙10-23 mol
Answer: 2.66 ∙ 10-8 cm and 1,992∙10-23 mol
Sources:
1. https://en.wikipedia.org/wiki/Kissing_number_problem
2. https://en.wikipedia.org/wiki/Amount_of_substance
3. https://en.wikipedia.org/wiki/Avogadro_constant